OpenStudy (anonymous):

help plz, sqrt-108 + sqrt-27 operations of complex numbers

OpenStudy (saifoo.khan):

15.5884573 i

OpenStudy (anonymous):

could you keep it in simplified form no decimal please

OpenStudy (saifoo.khan):

ok, i can try.

OpenStudy (saifoo.khan):

Myinin is here to save the Day!

myininaya (myininaya):

\[\sqrt{108}i+\sqrt{27}i=\sqrt{4}\sqrt{27}i+\sqrt{27}i=\sqrt{27}i(\sqrt{4}+1)=\sqrt{9}\sqrt{3}i(2+1)\] \[3i \sqrt{3}(3)=9i \sqrt{3}\]